# Fail Fast, Win Big: Rethinking the Drafting Strategy in Speculative Decoding via Diffusion LLMs

### v1

@@ Abstract
Diffusion Large Language Models (dLLMs) offer fast, parallel token generation, but their standalone use is plagued by an inherent efficiency-quality tradeoff. We show that, if carefully applied, the attributes of dLLMs can actually be a strength for drafters in speculative decoding with autoregressive (AR) verifiers. Our core insight is that dLLM’s speed from parallel decoding drastically lowers the risk of costly rejections, providing a practical mechanism to effectively realize the (elusive) lengthy drafts that lead to large speedups with speculative decoding. We present FailFast, a dLLM-based speculative decoding framework that realizes this approach by dynamically adapting its speculation length. It “fails fast” by spending minimal compute in hard-to-speculate regions to shrink speculation latency and “wins big” by aggressively extending draft lengths in easier regions to reduce verification latency (in many cases, speculating and accepting 70 tokens at a time!). Without any fine-tuning, FailFast delivers lossless acceleration of AR LLMs and achieves up to 4.9 speedup over vanilla decoding, 1.7 over the best naive dLLM drafter, and 1.4 over EAGLE-3 across diverse models and workloads. We open-source FailFast at https://github.com/ruipeterpan/failfast.
@@ 1 Introduction
Across diverse models and workloads, FailFast achieves a speedup of 1.9-4.9 over vanilla decoding, 1.2-1.7 over a strong dLLM drafter baseline, and 1.1-1.4 over EAGLE-3. FailFast achieves lossless acceleration of autoregressive target model generation while employing off-the-shelf dLLMs as draft models with no additional training/fine-tuning.
@@ 3.2 The Varying Decoding Difficulty within Sequences
In Easier Regions (e.g., tokens 150-380 for query 4): A short, fixed speculation length results in undue verification overhead. Even though the drafter could correctly guess the next 10 tokens, its token generation is forced to pause every 10 tokens to invoke the target model for verification. Since target model inference is memory-bound, the frequent loading of model weights incurs significant latency overhead, preventing the system from reaching peak overall efficiency.
@@ 5.1 Setup
EAGLE-3 (Li et al., 2025): EAGLE-3 is one of the most widely deployed state-of-the-art speculative decoding drafters. As a lightweight single-layer autoregressive Transformer, EAGLE-3 introduces a multi-layer feature fusion and a training-time test mechanism for improving drafting quality. Our EAGLE-3 drafter includes 1.38B parameters.
@@ 5.2 End-to-End Results
Comparisons with EAGLE-3. For Qwen2.5-7B, FailFast achieves a speedup of 1.1-1.4 over EAGLE-3 across datasets. Even though EAGLE-3 has a smaller drafting latency in each of its forward passes – it only has one layer compared to the 28 layers in other baselines, reducing the kernel launch overhead – it remains an autoregressive drafter and only generates a single token in each forward pass, whereas FailFast generates multiple tokens in each forward pass. Further, EAGLE-3 only drafts 5 tokens in each round, and is not able to capitalize on regions of easier tokens as a result. Even if a token tree with, for example, 64 tokens was used in EAGLE-3 drafting, the upper bound on accepted tokens would still be bottlenecked by the depth of the tree (up to 8 in EAGLE-3 (Li et al., 2025)).
@@ 6 Conclusion
Acknowledgments We thank Princeton’s Systems for Artificial Intelligence Lab (SAIL) and Princeton Language and Intelligence (PLI) for providing the hardware resources for running experiments. Rui would like to thank Minghao Yan for a random conversation on speculative decoding back in 2024 that planted the seed of this work.
@@ Appendix B Detailed Performance Comparison
| FailFast | Acceptance Rate | 68.9% | 66.8% | 61.9% | 61.3% | 73.9% | 66.6% |
@@ D.3 EAGLE-3
All EAGLE-3 baselines are run using vLLM v0.13.0 (Kwon et al., 2023). We use a num_speculative_tokens of 5, which was the speculation length that yielded the best speedup in our parameter sweeps. There are no public EAGLE-3 weights available for Qwen2.5. For Qwen2.5-7B-Instruct, we conducted pretraining of EAGLE-3 weights using the ultrachat-200K (Ding et al., 2023) text corpus in SpecForge (spe, 2025). We adopted the default training hyperparameters in SpecForge and trained EAGLE-3 to match the target model’s output. We are currently training our own EAGLE-3 weights on additional models and enhancing the current EAGLE-3 results. These comparison results will be included in a future version of the paper.

### v2

@@ Abstract
Diffusion Large Language Models (dLLMs) offer fast, parallel token generation, but their standalone use is plagued by an inherent efficiency-quality tradeoff. We show that, if carefully applied, the attributes of dLLMs can actually be a strength for drafters in speculative decoding with autoregressive (AR) verifiers. Our core insight is that dLLM’s speed from parallel decoding drastically lowers the risk of costly rejections, providing a practical mechanism to effectively realize the (elusive) lengthy drafts that lead to large speedups with speculative decoding. We present FailFast, a dLLM-based speculative decoding framework that realizes this approach by dynamically adapting its speculation length. It “fails fast” by spending minimal compute in hard-to-speculate regions to shrink speculation latency and “wins big” by aggressively extending draft lengths in easier regions to reduce verification latency (in many cases, speculating and accepting 70 tokens at a time!). Without any fine-tuning, FailFast delivers lossless acceleration of AR LLMs and achieves up to 4.9 speedup over vanilla decoding, 1.7 over the best naive dLLM drafter, and 2.0 over EAGLE-3 across diverse models and workloads. We open-source FailFast at https://github.com/ruipeterpan/failfast.
@@ 1 Introduction
Across diverse models and workloads, FailFast achieves a speedup of 1.9-4.9 over vanilla decoding, 1.2-1.7 over a strong dLLM drafter baseline, and 1.1-2.0 over EAGLE-3. FailFast achieves lossless acceleration of autoregressive target model generation while employing off-the-shelf dLLMs as draft models with no additional training/fine-tuning.
@@ 3.2 The Varying Decoding Difficulty within Sequences
In Easier Regions (e.g., tokens 150-380 for query 4): A short, fixed speculation length results in undue verification overhead. Even though the drafter could correctly speculate the next 10 tokens, its token generation is forced to pause every 10 tokens to invoke the target model for verification. Since target model inference is memory-bound, the frequent loading of model weights incurs significant latency overhead, preventing the system from reaching peak overall efficiency.
@@ 5.1 Setup
EAGLE-3 (Li et al., 2025): EAGLE-3 is one of the most widely deployed state-of-the-art speculative decoding drafters. As a lightweight single-layer autoregressive Transformer, EAGLE-3 introduces a multi-layer feature fusion and a training-time test mechanism for improving drafting quality. Similar to our other baselines, we also pick the best speculation length for each model-dataset combination.
@@ 5.2 End-to-End Results
Comparisons with EAGLE-3. Across models, FailFast achieves an average speedup of 1.2-1.7 over EAGLE-3 (up to 1.4-2.0 on specific datasets). Even though EAGLE-3 has a smaller drafting latency in each of its forward passes – it only has one layer compared to the 28 layers in other baselines, reducing the kernel launch overhead – it remains an autoregressive drafter and only generates a single token in each forward pass, whereas FailFast generates multiple tokens in each forward pass. Further, EAGLE-3 uses a fixed speculation length and is not able to capitalize on regions of easier tokens as a result. Even if a token tree with, for example, 64 tokens was used in EAGLE-3 drafting, the upper bound on accepted tokens would still be bottlenecked by the depth of the tree (up to 8 in EAGLE-3 (Li et al., 2025)).
@@ 6 Conclusion
Acknowledgments We thank Princeton’s Systems for Artificial Intelligence Lab (SAIL) and Princeton Language and Intelligence (PLI) for providing the hardware resources for running experiments. Rui would like to thank the many full-time researchers and interns at Google SRG for their help and support, especially Nikos Pagonas, Howie Chen, Manos Chatzakis, Vinay Banakar, Liana Patel, Jiani Zhang, and Lily Tsai. In addition, Rui would like to thank Minghao Yan for a random conversation on speculative decoding back in 2024 that planted the seed of this work.
@@ Appendix B Detailed Performance Comparison
| Fast-dLLM | Acceptance Rate | 68.9% | 66.8% | 61.9% | 61.3% | 73.9% | 66.6% |
@@ D.3 EAGLE-3
All EAGLE-3 baselines are run using vLLM v0.13.0 (Kwon et al., 2023). We run an extensive parameter sweep of num_speculative_tokens from 3 to 20, and pick a different speculation length that yielded the best speedup for each model-dataset pair. There are no public EAGLE-3 weights available for Qwen2.5. Thus, we conducted pretraining of EAGLE-3 weights using the ultrachat-200K (Ding et al., 2023) text corpus in SpecForge (spe, 2025). We adopted the default training hyperparameters in SpecForge and trained EAGLE-3 to match the target model’s output until convergence. We plot the moving average of the training accuracy in Fig. 7. During evaluation, we use the model checkpoints after five epochs of training (260k training steps). We find that even though further training improves training-time accuracy, they have a negligible impact on the end-to-end speedup of EAGLE-3. We open-source our EAGLE-3 model weights and training scripts.